OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-2, 3) and is perpendicular to 2x – 9y = -12 Please show all of your work.

OpenStudy (anonymous):

If it is perpendivular the slope is the opposite reciprocal of the slope of the line 2x-9y=-12 which has a slope of 2/9. Therefore the perpendicular line has a slope of -9/2 or -4.5 y=-4.5x+b using the point (-2,3) 3=-4.5(-2) + b b=-6 y=-4.5x-6

OpenStudy (anonymous):

why is this not better asked in mathematics?